OpenStudy (anonymous):

um what is equivalent to 3/10?

OpenStudy (anonymous):

there are an infinite number od equilvanet fractions. just multiple the numerator and denominator by the same number (top and bottom) 03*9 = 27 --- = --- 10*9 = 90 03*90 = 270 --- = ------ 10*90 = 900